OpenStudy (anonymous):

A 10-gallon container of water costs (y + 2) dollars while a 5-gallon container of water costs (y - 2) dollars. If the price per gallon of water is the same for both containers, the equation below can be used to solve for y. All of y plus 2 divided by 10 equals all of y minus 2 divided by 5 What is the value, in dollars, of y?

hero (hero):

\[\frac{y+2}{10} = \frac{y-2}{5}\] 5(y+2) = 10(y-2) 5y + 10 = 10y - 20 20 + 10 = 10 y - 5y 30 = 5y 30/5 = y 6 = y The value of y in dollars is $6.